OpenStudy (anonymous):

how to get from this t^2-2t-8 to this (t-4)(t+2)

OpenStudy (anonymous):

foil t*t=t^2 t*2=2t -4*t=-4t -4*2=-8 t^2+2t-4t-8 = t^2-2t -8

OpenStudy (anonymous):

I think she was asking how to factor, not how to multiply out the factors.

OpenStudy (anonymous):

yesss i was! i cant remember how to do in a simply way!

OpenStudy (anonymous):

when you have \[t^2-2t-8\] you're looking for factors like (t + a)(t + b), where a+b = -2 (from the 2t) and a x b is -8.

OpenStudy (anonymous):

sto factor....one method is to first look at the back number and fnd it;s factors

OpenStudy (anonymous):

so: what are two factors of -8, that when added result in -2?

OpenStudy (anonymous):

And then "-4 and +2" pops into your mind.

OpenStudy (anonymous):

so (t - 4)(t + 2)

OpenStudy (anonymous):

thank you! i am recalling all the algebra i learned in HS, thank u again!

OpenStudy (anonymous):

no prob!